OpenStudy (anonymous):

The Sun radiates away energy at the rate of 3.9*10^24 Watts. At what rate it it losing mass due to this radiation? E = Δmc^2 E = mc^2 - m0c^2 Where m0 rep. rest mass. For this one, I am unsure as to how I would go about with solving it. I know that a Watt can equal 1 Joule / second...

OpenStudy (anonymous):

Just as Energy = mass * c^2, Power = mass consumption rate * c^2 (Just divided either side with time) Therefore, mass consumption rate = Power / C^2 = 3.9 * 10^24 / (3*10^8)^2 = 43.33*10^6 kg/s